Operations Report – January 10th

Crew 202 Operations Report 10-JAN-2019

SOL: 12

Name of person filing report: Kasey Hilton

Non-nominal systems: Battery in suit #9

Notes on non-nominal systems: Battery in suit #9 was also very low after the EVA; Suit was checked to make sure the connections were proper and plugged in; I will check the voltage in morning once the battery has had time to charge

Generator (hours run): 17hr 52min; Turned on last night (09Jan2019) at 15:15; Turned off this morning (10Jan2019) at 9:07; Turned on tonight (10Jan2019) at 16:05

Solar SOC – Turned on (09Jan2019) 82%; Turned off (10Jan2019) 100%; Turned on (10Jan2019) 84%

Diesel Reading – 70%

Propane Reading – 35%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – 50%; 245 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 55%; 165 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 19:53 to refill tank

Water Meter: 01399491

Toilet tank emptied: Yes, stinky as always ://

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes

Hours: 66.8

Beginning charge: 100%

Ending charge: 72%

Currently charging: Yes

Opportunity rover used: No, brakes are still not working correctly

Hours: 45.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 69.0

Beginning charge: 100%

Ending charge: 89%

Currently charging: Yes

Notes on rovers: Brake fluid in Opportunity was replaced but the brakes are still not functioning properly

ATV’s Used: 350.1 (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: Needed an extra person on an EVA for film crew

Oil Added? No

ATV Fuel Used: Fuel is about one inch from the top after EVA

# Hours the ATVs were Used today: 1hr 46min

Notes on ATVs: Nothing to report

HabCar used and why, where? Not used

CrewCar used and why, where? Off site

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Battery in suit #7 had to be replaced again because the voltage was still very low after being sufficiently charged. Battery in suit #9 was low after the EVA as mentioned in “Non-nominal systems”; Suit will be checked after charging to see if the battery needs to be replaced or not

Summary of Hab operations: Hab ceiling still leaks every once and a while; Due to ice on the outside but should be gone in the next day or two

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: What are the average lifetime of a space suit battery? Suit #7 and #9’s batteries were just replaced a few days ago but are already below 12 V after a sufficient charge.